OpenStudy (anonymous):

User polar coordinates to evaluate the integral. double integral of (x^2) / (x^2+y^2) R is bounded by x^2+y^2=a^2 and x^2+y^2=b^2 0 < a < b

OpenStudy (anonymous):

how do you think we should start off this problem?

OpenStudy (anonymous):

i have absolutely no freaking idea LOL

OpenStudy (anonymous):

well you remeber those formulas right? x=rcos(theta) y=rsin(theta) x^2+y2=r^2

OpenStudy (anonymous):

so start by converting the function your are given to integrate into polar form, by replacing the x^2 with r^2cos^2(theta) and so on..

OpenStudy (anonymous):

i'm guessing convert fx into polars so r^2cos^2(theta) / r^2

OpenStudy (anonymous):

i'm assuming my theta region is 0 to theta, but i'm not quite sure what to do with r.. what's an annular region? lol

OpenStudy (experimentx):

a, b

OpenStudy (anonymous):

or it it should be the area lying between both circles

OpenStudy (anonymous):

so the limits for r should be from a to b and the limits for theat should be from 0 to 2pi

OpenStudy (experimentx):

Perhaps ... \[ \int_{0}^{2\pi}\int_{a}^{b} \cos^2 \theta\; r\;drd\theta\]

OpenStudy (anonymous):